# Novel View Synthesis for Cinematic Anatomy on Mobile and Immersive Displays

### v1

@@ Abstract
Interactive photorealistic visualization of 3D anatomy (i.e., Cinematic Anatomy) is used in medical education to explain the structure of the human body. It is currently restricted to frontal teaching scenarios, where the demonstrator needs a powerful GPU and high-speed access to a large storage device where the dataset is hosted. We demonstrate the use of novel view synthesis via compressed 3D Gaussian splatting to overcome this restriction and to enable students to perform cinematic anatomy on lightweight mobile devices and in virtual reality environments. We present an automatic approach for finding a set of images that captures all potentially seen structures in the data. By mixing closeup views with images from a distance, the splat representation can recover structures up to the voxel resolution. The use of Mip-Splatting enables smooth transitions when the focal length is increased. Even for GB datasets, the final renderable representation can usually be compressed to less than 70 MB, enabling interactive rendering on low-end devices using rasterization.
keywords:
Novel view synthesis, cinematic volume rendering, 3D Gaussian splatting
@@
Left: path-traced image of a 36 GB HiP-CT dataset, rendered in 95 seconds on a high-end GPU. Right: Gaussian splat representation of the same dataset requiring 69 MB and rendered at 60 frames per second. The rendering resolution is 2048x2048 pixels.
Introduction
Cinematic Anatomy is an immersive anatomy learning application developed by Siemens Healthineers, which is designed to improve anatomy education through the use of photorealistic 3D visualization and cinematic rendering[12]. Instead of real 3D anatomy models it utilizes volume data provided by medical devices such as computed tomography (CT) and magnetic resonance imaging (MRI) scanners.
The application is used in the field of anatomy education, specifically for teaching the diverse and complex individual human anatomy, anatomical variations, pathology, and age-related degeneration. It aims to enhance learners’ competency with immersive photorealistic 3D visualization of anatomy. In contrast to idealized anatomy models, learning with image data from real patients provides a complete experience of the patient journey from hospitalization to rehabilitation.
@@ Introduction
Additionally to stereoscopic projection modes for frontal education, Cinematic Anatomy also supports augmented and virtual reality headsets for personalized learning experiences. The portability of the created content is however often limited by the data size, especially when employing data from high-resolution imaging modalities like photon-counting CT, 7 Tesla MRI and phase-contrast CT[50]. Thus, Cinematic Anatomy is mostly used in frontal teaching scenarios, where the demonstrator uses a powerful GPU and has high-speed access to a large storage device where the dataset is stored.
We demonstrate that the limitations of Cinematic Anatomy can be addressed via novel view synthesis [35], a computer vision technique that reconstructs a volumetric radiance field of a scene from images of this scene. The reconstructed field can then be rendered with direct volume ray-casting from arbitrary views. With novel view synthesis from pre-rendered anatomy images, time consuming path-tracing can be avoided at demonstration time. On the other hand, visualization parameters such as transfer functions are baked into the volumetric radiance field and cannot be changed. We address this restriction by using compressed 3D Gaussian splatting for novel view synthesis. The resulting renderable representation has such a small memory footprint that multiple representations under different transfer function settings can be prepared and stored.
3D Gaussian splatting (3DGS) has been proposed by Zwicker et al. [61] to efficiently compute the projections of 3D Gaussian kernels onto the 2D image plane. Differentiable 3DGS [25] optimizes the number and parameters of the Gaussian kernels that are used to represent the scene. The Gaussians are optimized in a training process to reproduce initial images of the scene when rendered. Compressed 3DGS [38] has significantly reduced the memory consumption of the reconstructed Gaussian representation. Since it uses GPU rasterization, it runs efficiently even on mobile devices and can be integrated seamlessly into VR/AR environments. Novel View Synthesis for Cinematic Anatomy on Mobile and Immersive Displays demonstrates these properties with a high-resolution CT scan.
Contribution. To enable compressed 3DGS for Cinematic Anatomy, we make the following contributions:
  • We present an automatic approach for finding a set of images that captures all potentially seen structures under the current transfer function setting.
@@
  • We extend 3DGS[61] with differentiable alpha channel rendering to create background-free reconstructions.
  • We embed Mip-Splatting [59] to account for different levels of detail and enable smooth transitions when the focal length is increased.
  • We analyze the quality, performance and memory requirements with a number of high-resolution medical datasets that are rendered using a publically available Cinematic Anatomy tool.
@@
Our results demonstrate that even for multiple transfer function settings the overall memory requirement is significantly below of what is required by the initial dataset. For each single setting, the renderable representation is so small that it can be quickly downloaded over low-bandwidth channels and rendered on mobile devices. This enables Cinematic Anatomy with datasets that are initially so large that they do not even fit into the VRAM of high-end GPUs. Rendering performance is about two orders of magnitudes faster than optimized path-tracing, with almost no perceptible loss of image quality. This even facilitates the use of cinematic anatomy in mobile VR/AR environments with GPU multi-view rendering for stereoscopic displays.
Limitations. Besides the outlined strengths, 3DGS comes with the following limitations for Cinematic Anatomy: Firstly, lighting conditions are baked into the 3D Gaussian representation and cannot be changed during rendering. Secondly, the use of preset transfer functions hinders exploratory tasks, where domain experts might want to interactively perform changes of transfer function parameters. Thirdly, for highly transparent settings the quality of 3DGS currently decreases, requiring further research on tailored solutions to effectively cope with such scenarios.
@@ 1 Related Work
Differentiable 3D Gaussian splatting (3DGS) [25] builds upon elliptical weighted average (EWA) volume splatting [61] to efficiently compute the projections of 3D Gaussian kernels onto the 2D image plane. In addition, the number and parameters of the Gaussian kernels that are used to model the scene are optimized with differentiable rendering. Mip-Splatting [59] modifies 3DGS by integrating anti-aliasing with a 3D smoothing and a 2D Mip filter. It achieves improved quality of novel views at scales the Gaussian representation has not been optimized for. A number of approaches have concurrently proposed to convert the 3D Gaussian representations generated by 3DGS into a more compact form [38, 27]. These works have shown that for typical scenes the memory requirements of 3DGS can be brought below 50 MBytes, without any noticeable differences in the reconstructed images.
@@
3DGS falls into the category of methods for novel view synthesis, and it overcomes in particular the difficulties of voxel-based approaches [44, 15] to deal with sparsity. Even though approaches based on adaptive hash grids [37], tensor decomposition [9] or variants using dedicated compression schemes [29, 42] can effectively reduce the required memory, due to the use of volume ray-casting they require high-end GPUs to achieve reasonable rendering performance. The same limitation holds for differentiable volume rendering [55], which, similar in spirit to 3DGS, optimizes optical properties on a dense voxel grid using image-based loss functions.
An alternative approach for generating a compact volumetric representation is by means of Scene Representation Networks [34, 11, 40], i.e., fully-connected neural networks which have been introduced initially to encode a surface model as an implicit 3D function. Lu et al. [30] demonstrate the use of SRNs for volume data compression. By overfitting a network to a volume dataset, the network learns a compact latent space representation from which the initial dataset can be reconstructed. This, however, comes at the expense of subsequent network evaluations during rendering. This makes even the performance of GPU-friendly implementations [53] fall significantly below the performance of 3DGS.
While 3DGS uses rendered images of a volume dataset to build a renderable 3D object representation for novel view synthesis, light-field rendering [28, 18] aims at generating novel views by interpolation between given images. Light-field rendering has been used to accelerate novel-view synthesis for volume data [49, 5, 8], yet it is fair to say that it cannot compete with 3DGS with respect to quality and memory requirements.
To generate the images that are used by 3DGS, we use Cinematic Rendering, a volumetric Monte Carlo path tracing algorithm developed by Siemens Healthineers. It is integrated into the teaching application Cinematic Anatomy, which was developed in collaboration between the Johannes Kepler University Linz and Siemens Healthineers. A first prototype was installed at the Ars Electronica Center in Linz in 2015 in the Deep Space 8k digital experience space [14], a 16-by-9-meter wall projection space providing stereoscopic 8k resolution experiences to more than 100 visitors at a time. Since then, regular interactive visualizations of clinical volumetric DICOM data have been made available to the public. Furthermore, it has been incorporated into anatomy lectures for medical students and staff. In 2021, the JKU medSPACE, a lecture space for teaching anatomy, located at the Johannes Kepler University in Linz was opened. Developed and implemented by the Ars Electronica Futurelab, the JKU medSPACE shows anatomy in quadruple stereoscopic 3D 4K projection at 14×7 meters. It utilizes Cinematic Anatomy to provide virtual anatomy education to medical students and staff. Several studies have shown the benefits of using photo-realistic volumetric rendering of clinical volume data for teaching and understanding anatomy[3, 4, 43, 17].
@@
While our work builds upon 3DGS to interactively perform Cinematic Anatomy, other works have previously attempted to improve the performance of path tracing via image denoising [20, 22, 21], photon mapping [60], illumination caching [62] and adaptive temporal sampling [33]. Despite achieving remarkable performance gains at high quality, it is worth noting that these approaches cannot overcome the major limitation of interactive Cinematic Anatomy, i.e., it requires a rendering system with enormous memory resources to host high-resolution medical scans, as well as huge amounts of computing power to perform ray tracing with such data.
@@ 2 Cinematic anatomy pipeline
The different stages of the proposed pipeline for cinematic anatomy are shown in Fig. 1. It starts with reading a medical dataset, for which then one or multiple so-called presets are selected by the user. A preset includes the transfer function setting as well as material classifications and fixed clip planes that are used to reveal certain anatomical structures. For each preset, a set of views capturing all potentially seen structures in the data at varying resolution are computed (cf. Section 3.1). In this way, also structures which are not seen when generating images with camera positions on a surrounding sphere are recovered in the final object representation. These views are handed over to a physically-based renderer, i.e., a volumetric path tracer, which renders one image for every view using the corresponding preset (cf. Section 3.2). Once all images for a selected preset have been rendered, 3DGS is used to generate a set of 3D Gaussians with shape and appearance attributes so that their rendering matches the given images. Once the Gaussians are computed via differentiable rendering, they are compressed using sensitivity-aware vector quantization and entropy encoding (cf. Section 3.3). The final compressed 3DGS representation is rendered with WebGPU using GPU sorting and rasterization of projected 2D splats, with a pixel shader that evaluates and blends the 2D projections in image space.
Figure 1: Cinematic anatomy pipeline for a $1510\times 1706\times 1415$ HiP-CT dataset requiring 3.6 GB of memory. Numbers below each stage indicate the required memory at this stage and its computation times. 3D Gaussian splatting (3DGS) optimization uses 99 path traced images, and first generates the raw Gaussian representation in 48 minutes before it is compressed to 33 MB in 5 minutes.
3 Methods
@@ 3.1 View selection
Our approach iteratively generates a sequence of different viewpoints, so that every new viewpoint can recover as many as possible voxels which have not (or not sufficiently) been covered so far. While we use this approach in a fully automated manner in this work, in practice an initial set of views would be placed uniformly around the object and additional views would be recommended and accepted or denied by a domain expert.
Our approach shares similarities with techniques that automatically optimize visualization parameters to find a single best viewpoint, using image-based loss functions to guide an optimizer toward this optimum  [23, 48, 46, 10, 47, 58]. Viewpoint quality metrics for isosurface rendering were developed by Takahasi et al. [45] and Marsaglia et al. [31], based on applying entropy to either the rendered images or derived fields such as depth. Bordoloi et al. [6] introduce the use of per-voxel significance measures to quantify image entropy. Different viewpoint search algorithms using a given number of views on a sphere surrounding the volume were investigated in [32]. While these approaches first generate many possible viewpoints and select the best one using sampling, Weiss and Westermann [54] embed the optimization into the volume rendering process and let the optimizer converge to the best viewpoint.
Once a preset is selected, our method starts with computing an occupancy volume indicating for every voxel whether it contains material or not. A per-voxel visibility volume is initialized to zero. In the first phase, a set of camera positions (always looking at the center of the dataset) are generated on an ellipsoid entirely containing the dataset. The camera is moved toward or away from the center using a random offset. For each camera pose, for every non-empty voxel the transmittance along a view ray to this voxel is computed, and the per-voxel visibility is updated with the maximum of the current value and the transmittance.
The second phase generates additional views that cover as many as possible previously unseen structures, i.e., it selects camera poses that maximize the voxel visibility $gain$, which is defined as the increase of the visibility values by the transmittance values that are computed for a camera pose. We use Bayesian optimization [36, 16] for generating the new poses, which solves the optimization problem $\max_{\theta\in D}f(\theta)$ for a black box function $f:D\rightarrow\mathbb{R}$ with domain $D$. In our setting, $f$ is the function that maps the camera parameters to the voxel visibility gain. The goal of Bayesian optimization is to minimize the number of function evaluations. For this, a probabilistic (usually Gaussian) surrogate model and an acquisition function are applied. The former expresses Bayesian believe about the output of the objective function derived from prior evaluations, while the latter is used for selecting the next set of parameters for evaluation of the objective function. We make use of a publically available Python library [39] for performing the optimization process. For the acquisition function, the Upper Confidence Bound (UCB) [7] is used with parameter $\kappa=10$.
We iteratively generate additional camera poses by initializing the Bayesian optimizer with $K=128$ randomly sampled candidate camera poses in each iteration. A candidate camera pose is discarded and re-sampled if the bounding box of the data lies outside the view frustum or the camera position is inside an occupied voxel. Then, Bayesian optimization is started to iteratively sample another set of $K$ candidate camera poses. Here, it utilizes the prior that has been captured through previous camera poses to guide the optimization process. Only the best candidate pose is finally selected per iteration.
@@ 3.2 Image generation
We render volume data with Monte Carlo volume path tracing  [12, 13, 2, 26, 24, 2] from multiple views to generate a set of training images. Rendering starts by tracing paths from the detector pixels of a virtual camera into the scene. A transfer function is used to map the scalar quantities of the 3D scans to an emissive color and density. Delta tracking [56] is applied to decide whether at a sample point a scattering event, absorption event or null collision occurred. In the case of an absorption event, the path is terminated and the emissive color is regarded as the contribution of the path. In the case of a scattering event, the Henyey-Greenstein phase-function [19] is applied to determine the next direction along which the ray is followed. In the case of a null-collision, the path is followed unchanged.
Additionally to the semi-transparent volumetric component, opaque isosurfaces can be embedded into the rendering process. For this, the local density gradient magnitude at each sampling point is computed, and an intersection with a surface is assumed when the gradient magnitude exceeds a user-specified iso value. The iso surface can also be colored by applying a transfer function that maps density to color. Global shading is computed by generating a reflection event for the surface with the new ray direction being sampled proportional to the probability density function (PDF) of a chosen bidirectional reflectance distribution function (BRDF) of the surface. This process is repeated until the ray leaves the volume domain or an absorption event happens. Absorption may be determined by a fixed maximum number of scatter events or using Russian roulette.
High dynamic range light maps are employed to look up lighting information contributions from the environment. We also support synthetic local light sources. Next event estimation is used to importance sample rays towards the light source and potentially reduce the variance of the rendered image. All Monte Carlo samples are accumulated and averaged in a floating-point accumulation buffer. A tone-mapping pass maps the accumulated result into the final lower dynamic range output buffer. For fast image generation, we apply performance optimization methods such as empty-space skipping and memory coherent scattering. The latter optimization ensures that rays of neighboring pixels are scattered in the same direction, thus ensuring optimized cache utilization.

### v2

# Application of 3D Gaussian Splatting for Cinematic Anatomy on Consumer Class Devices

@@ Abstract
Interactive photorealistic rendering of 3D anatomy is used in medical education to explain the structure of the human body. It is currently restricted to frontal teaching scenarios, where even with a powerful GPU and high-speed access to a large storage device where the data set is hosted, interactive demonstrations can hardly be achieved. We present the use of novel view synthesis via compressed 3D Gaussian Splatting (3DGS) to overcome this restriction, and to even enable students to perform cinematic anatomy on lightweight and mobile devices. Our proposed pipeline first finds a set of camera poses that captures all potentially seen structures in the data. High-quality images are then generated with path tracing and converted into a compact 3DGS representation, consuming < 70 MB even for data sets of multiple GBs. This allows for real-time photorealistic novel view synthesis that recovers structures up to the voxel resolution and is almost indistinguishable from the path-traced images.
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010372</concept_id> <concept_desc>Computing methodologies Rendering</concept_desc> <concept_significance>100</concept_significance> </concept> </ccs2012>
@@
[300]Computing methodologies Computer graphics; Rendering
[…]
1 Introduction
Cinematic Anatomy (CA) is an immersive anatomy learning application, which is designed to improve anatomy education through the use of photorealistic 3D rendering via path-tracing[CEGM16]. Instead of real 3D anatomy models, it utilizes volume data provided by medical scanning devices. The application is used in the field of anatomy education, e.g., in the JKU medSPACE [JKU], a lecture space for teaching anatomy. It is used for teaching the diverse and complex individual human anatomy, anatomical variations, and pathology, to enhance learners’ competency with immersive photorealistic 3D visualization of data from real patients. Several studies have shown the benefits of using photo-realistic volumetric rendering of clinical volume data for teaching and understanding anatomy [GES18, BKE19, SWS22].
@@ 1 Introduction
Additionally to stereoscopic projection modes for frontal education, students need to run CA on their mobile devices for personalized learning experiences. In addition to significant performance losses on such devices, the portability of the created content is however often limited by the data size, especially when employing data from high-resolution imaging modalities like photon-counting CT, 7 Tesla MRI and phase-contrast CT [WTW21]. Thus, CA is mostly used in frontal teaching scenarios, where the demonstrator uses a powerful GPU and has high-speed access to a large storage device where the data set is stored. As Application of 3D Gaussian Splatting for Cinematic Anatomy on Consumer Class Devices demonstrates, even then is it difficult to render the data at interactive rates.
We demonstrate that differentiable 3DGS [KKLD23], which reconstructs a 3D Gaussian scene representation from images of this scene, can address the limitations of CA. The Gaussian representation can be rendered at high speed from arbitrary views, avoiding time-consuming path tracing. In combination with compressed 3DGS [NSW24], the memory consumption of the Gaussian representation is significantly reduced, and with GPU rasterization, 3D Gaussian splatting runs efficiently even on mobile devices. Application of 3D Gaussian Splatting for Cinematic Anatomy on Consumer Class Devices demonstrates these properties with a high-resolution CT scan.
Contribution. To use compressed 3DGS for CA, we propose a processing pipeline including the following adaptations:
  • We extend the view selection method proposed by Kopanas and Drettakis [KD23] for volume rendering to automatically find a set of cameras that captures all potentially seen structures under the current transfer function setting.
@@
  • We extend 3DGS with differentiable alpha channel rendering to create background-free reconstructions and drastically improve the reconstruction of translucent materials.
@@
We analyze the quality, performance, and memory requirements with several high-resolution data sets. Training images are rendered with a publically available CA tool. The results demonstrate that the memory requirement is significantly below the initial data size. Since the renderable representation is so small, students can quickly download it over low-bandwidth channels and render on their mobile devices. Rendering performance is about two orders of magnitudes faster than optimized path tracing, with almost no perceptible loss of image quality.
Limitations. The use of 3DGS for CA comes with the following limitations: Firstly, lighting conditions are baked into the 3D Gaussian representation and cannot be changed during rendering. Secondly, due to the use of preset transfer functions and clip planes, the approach is less effective in supporting interactive volume exploration. Overcoming these limitations is difficult, and we discuss possible improvement strategies at the end of our work.
Figure 1: CA pipeline for a $1510\times 1706\times 1415$ HiP-CT data set requiring 3.6 GB of memory. Numbers below each stage indicate the required memory at this stage and its computation times. 3D Gaussian splatting (3DGS) optimization uses 99 path traced images, and first generates the raw Gaussian representation in 48 minutes before it is compressed to 33 MB in 5 minutes.
@@ 2 Related Work
3DGS [KKLD23] builds upon elliptical weighted average (EWA) volume splatting [ZPVBG01] to efficiently compute the projections of 3D Gaussian kernels onto the 2D image plane. In addition, the number and parameters of the Gaussian kernels that are used to model the scene are optimized with differentiable rendering. Mip-Splatting [YCH23] modifies 3DGS by integrating anti-aliasing with a 3D smoothing and a 2D Mip filter. It achieves improved quality of novel views at scales the Gaussian representation has not been optimized for. A number of approaches have concurrently proposed to convert the 3D Gaussian representations generated by 3DGS into a more compact form [NSW24, LRS24]. For typical scenes, the memory requirements of 3DGS is below 50 MB without any noticeable differences in the reconstructed images.
@@
3DGS for novel view synthesis overcomes in particular the difficulties of voxel-based approaches [MST20, FKYT22] to deal with sparsity. Even though adaptive hash grids [MESK22], tensor decomposition [CXG22] or variants using dedicated compression schemes [LSW23, RLN23] can effectively reduce the required memory, they use volume ray-casting and, thus, require high-end GPUs to achieve reasonable rendering performance. The same limitation holds for differentiable volume rendering [WW22], which, similar in spirit to 3DGS, optimizes optical properties on a dense voxel grid using image-based loss functions.
@@
3DGS optimizes a 3D scene representation from photorealistic images of that scene, which are generated with volumetric path-tracing [PJH23, NSJ14, NGHJ18, NDSRJ20]. A number of approaches have previously attempted to improve the performance of path tracing via image denoising [HMES20, JLM23, IGMM22], photon mapping [YYS23], illumination caching [vLMB23] and adaptive temporal sampling [MHK19]. Despite achieving remarkable performance gains at high quality, all these approaches require a rendering system with enormous memory resources to host high-resolution data sets as well as huge computational power to perform ray tracing with such data. It is fair to say that high-quality path-tracing on consumer class hardware is impossible today.
In principle, the memory requirements of CA can be addressed with Scene Representation Networks (SRNs)  [MON19, CZ19, PFS19], i.e., fully-connected neural networks that learn to encode a surface model as an implicit 3D function. Lu et al. [LJLB21] demonstrate the use of SRNs for volume data compression, by overfitting a network to a volume data set. This approach, however, comes at the expense of subsequent network evaluations during rendering, which makes even GPU-friendly ray-marching implementations [WHW22] significantly slower than 3DGS.
A challenging problem in novel view synthesis is to find an as small as possible set of camera poses for generating the training and test images. Note that this problem is different to the problem of viewpoint optimization in visualization, where visualization parameters are optimized to find a single best viewpoint, e.g., by using entropy-based  [JS06, VMN08, TLB09, CJ10, WHW22] or similarity-based [TWC16, YLLY19] loss functions that guide an optimizer. For SRNs, Kopanas and Drettakis [KD23] have introduced an algorithm to automatically optimize the placement of cameras so that improved coverage of a scene is achieved. We adapt this algorithm to work with volumetric data sets.
@@ 3 Cinematic Anatomy Pipeline
The different stages of the proposed CA pipeline are shown in Fig. 1. After loading a data set, one or multiple so-called presets are selected by the user. A preset includes the transfer function setting as well as material classifications and fixed clip planes that are used to reveal certain anatomical structures.
For each preset, multiple views capturing all potentially seen structures in the data are computed (cf. Section 3.1). In this way, we recover structures in the final object representation which are not seen when generating images with camera positions on a surrounding sphere. These views are handed over to a physically-based renderer, i.e., a volumetric path tracer, which renders one image for every view using the corresponding preset (cf. Section 3.2).
Once the images for a selected preset have been rendered, 3DGS generates a set of 3D Gaussian splats with shape and appearance attributes so that their rendering matches the given images. Once the parameters of the Gaussians are computed via differentiable rendering, they are compressed using sensitivity-aware vector quantization and entropy encoding (cf. Section 3.3). The final compressed 3DGS representation is rendered with WebGPU using GPU sorting and rasterization of projected 2D splats, with a pixel shader that evaluates and blends the 2D projections in image space. We embed Mip-Splatting [YCH23] to account for different levels of detail and enable smooth transitions when the focal length is increased.
@@ 3.1 View Selection
Novel view synthesis requires that all visible scene parts are covered in the training images. Kopanas and Drettakis [KD23] propose an automatic camera placement algorithm for SRNs which maximises observation frequency and angular uniformity. The observation frequency lies between 0 (no camera observes a point) and 1 (all cameras observe a point). The angular uniformity considers the total variation distance between a 2D histogram of the directions of cameras observing a point in spherical coordinates and a uniform distribution. The algorithm iterates over batches of 1000 randomly sampled camera poses. In each iteration the cameras lying inside of or too close to occupied space are rejected. From the remaining camera poses the one is selected resulting in the highest improvement of the reconstruction, measured by observation frequency and angular uniformity.
We propose two modifications of this algorithm for 3DGS of volumetric data sets. Firstly, we observe that due to the high dimensionality of the search space, a huge number of cameras needs to be evaluated and optimal camera poses might be missed. To avoid this, we use Bayesian Optimal Sampling (BOS) [Moc89, Gar23] to adaptively place cameras in regions that are more promising to yield an improved maximum (called exploitation) or in previously rather unexplored regions (called exploration). In this way, the chance of missing optimal camera poses is significantly reduced, and fewer training images need to be generated. Secondly, instead of using a binary visibility indicating whether a point lies inside or outside the camera frustum, we use a continuous visibility that also considers (partial) occlusion. At each voxel, GPU ray marching is performed to compute the maximum transmittance, which is then used as a visibility indicator. For data sets not fitting into GPU memory, we perform this step with a lower-resolution copy.
BOS applies a probabilistic (usually Gaussian) surrogate model and an acquisition function. The former expresses Bayesian belief about the output of the objective function derived from prior evaluations, and the latter is used for selecting the next set of parameters for evaluating the objective function. For the acquisition function, the upper confidence bound [BCdF10] is used with parameter $\kappa=10$, which controls the trade-off between exploitation and exploration. The value was empirically determined to work well with all test data sets, but can also be subjected to hyperparameter optimization either regarding the energy term by Kopanas and Drettakis [KD23] or the reconstruction quality with respect to images in a training set. We make use of the publically available software library Limbo [CCAM18] to perform the optimization process.
In Fig. 2, we showcase the ability of automatic view selection using BOS for improved capturing of internal structures of a concave 2D test data set. When randomly sampling cameras on the hemisphere around an object, structures in the interior are missed. When sampling only a small set of cameras on the hemisphere, BOS selects few additional camera poses which capture insufficiently seen structures on the outside and not yet captured structures on the inside. We demonstrate the resulting improvements in the reconstruction quality with 3DGS in Section 4.3. We further compare the convergence rate and performance of BOS and random sampling [KD23] in the supplementary material.
(a) Pure hemisphere sampling.
(b) Bayesian optimal sampling.
Figure 2: View selection strategies for a concave 2D test data set (black). Left: Random sampling of camera poses on a bounding circle cannot capture interior structures. Right: With BOS, given a sparse set of initial poses (grey), additional poses (red) are automatically determined to capture otherwise not seen object parts. Red area indicates seen domain parts.
@@ 3.2 Image Generation
We render volume data with Monte Carlo volume path tracing from multiple views to generate a set of training images. Delta tracking [WMHL65] is used to determine a scattering event (the Henyey-Greenstein phase-function [HG41] is applied to determine the next ray direction), an absorption event (the path is terminated and the emissive color is regarded as the path contribution) or a null collision (the path is followed unchanged). A surface intersection is assumed when the density gradient magnitude exceeds a user-specified iso-value. Global illumination is then simulated by generating a reflection event for the surface with the new ray direction sampled proportional to the probability density function of a chosen reflectance distribution function. This process is repeated until the ray leaves the volume domain or an absorption event happens.
High dynamic range light maps are employed to look up lighting information from the environment. Next event estimation is used to importance-sample rays towards the light source and potentially reduce the variance of the rendered image. All Monte Carlo samples are accumulated and averaged in a floating-point accumulation buffer. A tone-mapping pass maps the accumulated result into the final lower dynamic range output buffer. For fast image generation, we apply performance optimization methods such as empty-space skipping (based on the transfer function preset) and memory coherent scattering. The latter optimization ensures that rays of neighboring pixels are scattered in the same direction, thus ensuring optimized cache utilization.